We deal from a well shuffled 52-card deck. How many different ways can you deal 13 cards so that there are no king cards among the first 12?
Answer #1

The probability that the first card is not a king is 48/52. Given that, the probability that the second is not a king is 47/51. We continue similarly until the 12th card. The probability that the 12th card is not a king, given that none of the preceding 11 was a king, is 37/41. (There are 52 − 11 = 41 cards left, and 48 − 11 = 37 of them are not kings.) Finally, the conditional probability that the 13th card is a king is 4/40. The desired probability is
(48·47···37·4 )/ (52 · 51 · · · 41 · 40) = 0.0337575